OpenStudy (anonymous):

WHAT IS THE SUM FOR WHICH 2 YRS SIMPLE INTEREST AND COMPOUND INTEREST DIFFERENCE IS 880 , RATE 4 PER CENT ANN FOR BOTH

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

compounded how frequently?

OpenStudy (anonymous):

i make it 550,000

OpenStudy (anonymous):

let the sum be x. x(1.04)^2 - x*8 / 100=880 1.0816x - 0.08x = 880 1.0016x = 880 x = $878.60

OpenStudy (anonymous):

im not sure though

OpenStudy (anonymous):

s1 = p +p x .04 x 2 = 1.08p s2 = p(1 + r)^2 = 1.04^2 = 1.0816 s2-s1 = 0.0016 = 880 p = 880 / .0016 = 550000

OpenStudy (anonymous):

yeah i added the 1 by mistake...